Compute the monthly payment and the total amount spent for a vehicle that costs $15,500 if you finance the entire purchase over 5 years at an annual rate of 8.50 percent. Calculate the payment if you finance the car for only four years. Finally, calculate the payment for three years. What do you notice about the payment under the different time assumptions? Click on the table icon to view the MILPF table EEB The monthly payment, PMT, on the 5-year auto loan is s (Round to the nearest cent) The total amount spent if you financed $15,500 for 5 years at 8.50 percent per year is S(Round to the nearest cent.) The monthly payment, PMT, on the 4-year auto loan is (Round to the nearest cent.) Finally, the monthly payment, PMT, on the 3-year auto loan is (Round to the nearest cent.) What do you notice about the payment under the different time assumptions? (Select the best choice below.) 0 A. As the number of payments decreases, the monthly payment increases, all else equal. The shorter the borrowing period, the higher the monthly payments O B. As the number of payments decreases, the monthly payment decreases, all else equal. The shorter the borrowing period, the lower the monthly payments
Calculate how much money a prospective homeowner would need for closing costs on a house that costs $131,000. Calculate based on a 17 percent down payment, 1.4 discount points on the loan, a 0.7 point origination fee, and $1,270 in other fees. The closing costs would be s (Round to the nearest dollar.)
Annie's mortgage statement shows a total payment of $659.41 with $589.11 paid toward principal and interest and $70.30 paid for taxes and insurance. Taxes and insurance for three months were collected at closing. Now, after six months of payments, she is curious about the total in her escrow account. Calculate the amount for, her and explain the account. Calculate the amount for her and explain the account. (Select the best answer below.) 0 A. Assuming no monthly, quarterly or semi-annual tax or insurance withdrawals from the account, the account should total $210.90. An escrow account is a O B. Assuming no monthly, quarterly or semi-annual tax or insurance withdrawals from the account, the account should total $632.70. An escrow account is a ○ C Assuming no monthly, quarterly or semi-annual tax or insurance withdrawals from the account, he account should total $421.80. An escrow account is a special reserve account used to accumulate the annual property (real estate) tax payments and homeowners insurance premiums for the homeowner. special reserve account used to accumulate the annual property (real estate) tax payments and homeowner's insurance premiums for the homeowner special reserve account used to accumulate the annual property (real estate) tax payments and homeowner's insurance premiums for the homeowner.
Determine the maximum 30-year fixed-rate mortgage amount for which a couple could qualify if the rate is 6.14 percent. Assume they have other debt payments totaling $348 per month and a combined annual income of $71,100. Monthly escrow payments for real estate taxes and homeowner's insurance are estimated to be $284. (Assume a 36 percent maximum of annual income for total debt and escrow payments.) The maximum 30-year fixed-rate mortgage amount for which a couple could qualify if the rate s 6 14% is $ Round to the nearest dollar.
Answer #1

(1) Borrowing = $ 15500, Tenure = 5 years or 60 months and Interest Rate = 8.5 % per annum

Applicable Monthly Rate = 8.5 / 12 = 0.7083 %

Let the monthly payment be P1

Therefore, 15500 = P1 x (1/0.007083) x [1-{1/(1.007083)^(60)}]

15500 = P1 x 48.74164

P1 = 15500 / 48.74164 ~ $ 318

Total Amount Paid = 318 x 60 = $ 19080.2

New Tenure = 4 years or 48 months, Let the monthly payments be P2

Therefore, 15500 = P2 x (1/0.007083) x [1-{1/(1.007083)^(48)}]

15500 = P2 x 40.57105

P2 = 15500 / 40.5705 = $ 382.0458 ~ $ 382.05

Total Amount Paid = 382.05 x 48 = $ 18338.2

New Tenure = 3 years or 36 months, Let the monthly payments be P3

Therefore, 15500 = P3 x (1/0.007083) x [1-{1/(1.007083)^(36)}]

15500 = P3 x 31.6783

P3 = 15500 / 31.6783 = $ 489.2939 ~ $ 489.29

Total Amount Paid = 489.29 x 36 = $ 17614.44

As the number of payments decrease, the monthly payments increase.

Hence, the correct option is (A)

NOTE: Please raise separate qeuries for solutions to the remaining unrelated questions, as one query is restricted to the solution of only one complete question with upto 4 sub-parts.